# Truly Assessing Fluid Intelligence of Large Language Models through Dynamic Reasoning Evaluation

### v1

@@ 4.1 Experimental Settings
Evaluated LLMs. For completeness, we test 11 representative LLMs varying in parameters, vision encoders, including close-sourced APIs and open-sourced LLMs. Close-sourced APIs from different companies encompass GPT-4o openai2024gpt4o , OpenAI-o1 openai2024gpto1 , Claude-3.7 2024claude and OpenAI-o3-mini openai2025o3mini . Open-sourced LLMs include DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR , QwQ, Qwen2.5 yang2024qwen2 , and Skywork-OR1 skywork-or1-2025 . See Supplementary Materials for details of evaluated LLMs. To reduce randomness, all presented results of models are average results over three trials.
Evaluation Methods. In the DRE-Bench benchmark, accuracy serves as the primary evaluation metric, defined as the proportion of samples for which the model’s output grid exactly matches the ground-truth output grid. To avoid contingency, each variable contains an average of 11 samples for each of its values. All inferences are performed using the vLLM backend kwon2023efficient . To ensure fairness and consistency, we adopt the official standardized prompting template released by ARCPrize arcprize2024modelbaseline . We conducted all inference evaluations of open-source models on 8 NVIDIA A100 (80G) GPUs, while performing evaluations of APIs on CPUs.
4.2 Main Results in Four Levels
Furthermore, as task difficulty increases, performance disparities among models become more pronounced, highlighting the potential of incorporating dedicated reasoning paradigms for addressing fluid intelligence problems. For Level 3B tasks, which require conceptual knowledge, all existing models fail, underscoring the current limitations of even advanced reasoning models. These findings emphasize both the inherent challenges posed by our benchmark and its flexibility in revealing model capabilities across a wide spectrum of cognitive demands.
@@ 4.3 Dynamic Trends across Different Cognitive Levels
Since our generator is capable of producing data with varying levels of complexity, we conduct a fine-grained evaluation to assess model performance across different cognitive levels. Figure 14 illustrates representative performance curves corresponding to each cognitive level. Additional results are provided in Supplementary Materials.
In Level-1, models exhibited relatively uniform performance patterns on the count task. Since these tasks involve basic enumeration without substantial cognitive demands, the models consistently achieved high average accuracy, and increases in complexity had minimal impact.
For Level-2, performance gaps among models became increasingly pronounced. Even as the complexity of these spatial tasks decreased, lower-performing models continued to struggle considerably. Impressively, models that excelled in overall evaluation remained robust and relatively unaffected by complexity increase, maintaining stable accuracy with minor fluctuations. This observation suggests that these models have, to some extent, acquired the capability to resolve spatial reasoning problems.
Regarding tasks in Level-3, we observe a substantial performance drop as the number of required planning steps increases. Most models can only manage the simplest scenarios, with a consistent failure point emerging when the planning depth reaches two steps. This highlights substantial limitations in the current intelligence of LLMs, suggesting that they have not genuinely mastered these planning-based rules.
Finally, at Level-4, even when the complexity of the cases is reduced, models still fail to provide correct solutions. This finding indicates that current models have only a rudimentary grasp of physical concepts and have yet to internalize even the most fundamental principles of intuitive physics.
@@
To further illustrate the performance and stability of each model on dynamic tasks, Figure 5 presents the mean accuracy and corresponding variance across different cognitive levels. As shown in the figure, for the majority of Level-1 tasks, OpenAI-o1, DeepSeek-R1, and Claude-3.7 demonstrate strong performance and high stability. However, when the task difficulty increases to Level-2, Claude-3.7 exhibits substantial fluctuations in performance, indicating limited generalization capabilities at this level. In contrast, OpenAI-o1 and DeepSeek-R1 maintain comparable performance and stability to those observed at Level-1, highlighting the advantage of reaoning models in solving more cognitively demanding tasks. Moreover, in the third panel of Figure 5, most of the scatter points are concentrated in the lower-left region, suggesting that current models struggle to generalize effectively across the more complex and varied tasks at Level-3.
@@ 4.4 Ablation Study
Impact of the Number of In-context Learning Samples. Previous work brown2020language ; openai2023gpt4 has demonstrated the effectiveness of in-context learning in enhancing model performance across a variety of tasks. In this section, we vary the number of training examples to investigate how the quantity of in-context samples affects performance in our specific scenario. The results are shown in Figure 7. Overall, increasing the number of in-context samples helps models better capture underlying rules and improve performance. In higher levels like Level-2, Level-3 and Level-4, increasing the number of in-context training samples leads to noticeable performance improvements. However, for Level-1 tasks, increasing the number of samples yields limited improvement. This suggests that adding more in-context examples has limited impact when the model has already mastered the task or lacks the inherent capability to solve it. In both cases, the effectiveness of in-context learning is diminished.
Impact of the Auxiliary Visual Information. Previous studies legris2024h ; patterson2014human have shown that humans tend to perform better on abstract reasoning tasks when the grids are visualized, as visualization can aid in recognizing patterns and rules. Motivated by these findings, we investigate whether adding auxiliary visual information can enhance model performance. Specifically, we visualize each case and experiment with two formats: single-image, which presents all three training input-output pairs along with the test input in a single image; multi-image, which provides them as seven separate images. Table 7 presents the experimental results of GPT-4o and Claude 3.7 across cases spanning all four cognitive reasoning levels. Overall, neither visual format consistently outperforms the text-only baseline, and in some instances, accuracy even declines. These results suggest that current models struggle to derive meaningful improvements in abstract reasoning from visualized inputs.
Figure 6: Accuracy of DeepSeek-R1 on different numbers of in-context training samples.
Figure 7: Accuracy comparison between text-only(-), single-image(single), and multi-image(multi) settings, across four cognitive reasoning levels.
@@
Single 96.84 17.14 31.25 15.87
Multi 97.89 17.14 35.00 12.70
Impact of the Inference Time. Recent studies DeepSeekAI2025DeepSeekR1IR ; openai2024gpto1 ; qin2024o1 ; huang2024o1 have demonstrated that inference-time scaling plays a crucial role in enhancing model performance on complex tasks. Building upon these, we take a step to examine how the model’s inference time varies with increasing task complexity. Since the internal reasoning process of O1 is inaccessible, we use the response latency to measure for the inference time. The results are presented in Figure 9. We observe that as task complexity increases, the model tends to engage in deeper reasoning. However, in tasks at Level-3 (i.e., planning), where the model can only solve problems of relatively low complexity, simply increasing inference time is insufficient to compensate for the model’s inherent limitations.
@@
Figure 9: Results of directional move and symmetry.
@@ 4.5 Case Study
Analysis of Spatial Orientations. Upon closer examination of the results, we find that current models may demonstrate a distinct understanding of spatial orientation compared to humans. As shown in Table 9, the models achieve higher and more consistent accuracy in vertical (up/down) directions than in horizontal (left/right) ones in Move. Similarly, in symmetry tasks, performance is better for horizontal symmetry than for vertical symmetry. However, from the perspective of human cognition, directional distinctions are typically perceived as equivalent aflalo2008four ; ambinder2009human . These findings suggest that current LLMs may exhibit systematic divergences from human cognitive patterns in processing spatial orientation.
@@
Analysis of Error Cases. As shown in Figure 10, we randomly select error cases from four cognitive levels where we ensure the cases are of comparable complexity. Then we visualize them alongside the corresponding ground-truth outputs for comparison. In Level-1 and Level-2, the differences between the model’s incorrect predictions and the correct answers are relatively subtle—indicating that the model roughly understands the required operation. However, in Levels-3 and Level-4, the incorrect outputs become significantly more disorganized and divergent from the ground truth, suggesting a complete failure to grasp the underlying rule. This is especially evident in Level-4, where physical concepts pose substantial challenges to the models. These observations highlight that as the cognitive level increases, the nature of model errors becomes increasingly complex and unreasonable.
@@
Figure 10: Error cases on OpenAI-o1 across four cognitive levels. For each error case example, we visualize the input, together with the ground truth and model output.
5 Conclusion and Limitation
@@
Conclusion In this work, we present DRE-Bench, a benchmark designed to evaluate the fluid intelligence of large language models (LLMs) through abstract reasoning tasks grounded in a structured four-level cognitive hierarchy. Unlike prior benchmarks, DRE-Bench offers interpretability and dynamic robustness by probing latent rules across multiple tasks and variants at each cognitive level. Our results show that while LLMs have advanced, fluid intelligence remains elusive. DRE-Bench offers a scalable and interpretable framework for analyzing and tracking reasoning capabilities, paving the way for future models with stronger generalizable reasoning.
Limitation. Our dynamic reasoning benchmark is fully code-generated, ensuring 100% accuracy and verifiability. While this approach guarantees data quality, it requires annotators to manually implement task-specific generators and solvers, making task expansion labor-intensive. Moreover, as discussed in Section 4.5, incorrect model responses exhibit varying degrees of deviation, suggesting that accuracy alone may be inadequate and finer-grained evaluation metrics are needed to better assess reasoning quality.
Appendix A Overview of Appendix
Due to the page limitation of the manuscript, we provide more information and details about our benchmark in the following aspects:
  • Sec. B: Braoder Impact.
  • Sec. C: Details of DRE-Bench.
    • Sec. C.1: Detailed Dataset Content and Distribution.
    • Sec. C.2: Dataset Document.
  • Sec. D: Experimental Details.
    • Sec. D.1: Details of Evaluated LLMs.
    • Sec. D.2: More Dynamic Evaluation Curves.
    • Sec. D.3: Visualization Format.
    • Sec. D.3: Table of Accuracy and Stability.
Appendix B Broader Impacts
The development of DRE-Bench represents a meaningful step toward more principled, interpretable, and verifiable evaluation of fluid intelligence in large language models. By structuring reasoning tasks within a cognitively grounded hierarchy and dynamically controlling rule-related complexity, DRE-Bench enables fine-grained assessment of models’ genuine reasoning abilities rather than surface-level pattern matching or memorization.
Moreover, the benchmark emphasizes a scalable and reproducible methodology: all task samples and solutions are generated deterministically via code, ensuring full verifiability of the evaluation process. This offers a transparent framework not only for performance comparison but also for diagnosing specific weaknesses in LLM reasoning across cognitive levels. As such, DRE-Bench can serve as a foundation for future work in designing cognitively aligned models with deeper, more structured reasoning processes.
Nonetheless, optimizing solely for DRE-Bench performance may result in artificial improvements that do not reflect true reasoning competence. Therefore, we encourage the community to treat DRE-Bench as one component within a broader suite of diverse cognitive evaluations. Future research should also explore how the cognitive insights derived from structured reasoning benchmarks can inform more human-aligned training objectives, leading to AI systems that are not only more capable, but also more interpretable, robust, and trustworthy in high-stakes environments.
@@ C.1 Detailed Dataset Content and Distribution
Symmetry Perform {Vertical, Horizontal, Center} symmetry of the object. number {1, 8} 75
Sequential Categorization Classify objects based on examples, and apply the corresponding rule to each category. category {1, 5} 65
@@
Conceptual Gravity Objects in mid-air should fall downward according to gravity. number {1, 8} 63
Reflection The light reflects upon hitting walls. number {1, 8} 100
Expansion Objects expand when heated until obstructed. number {1–8} 50
@@ C.2 Dataset Document
We provide comprehensive documentation of our dataset along with its intended use cases. The dataset and accompanying resources are available at the following link: https://huggingface.co/datasets/yang5114/DRE-Bench, which includes metadata, format details, and so on.
Appendix D Experimental Details
D.1 Details of Evaluated LLMs

### v2

@@ 4.1 Experimental Settings
Evaluated LLMs. For completeness, we test 11 representative LLMs varying in parameters, vision encoders, including close-sourced APIs and open-sourced LLMs. Close-sourced APIs from different companies encompass GPT-4o (OpenAI, 2024a), OpenAI-o1 (OpenAI, 2024b), Claude-3.7 (Anthropic, 2024) and OpenAI-o3-mini (OpenAI, 2025). Open-sourced LLMs include DeepSeek-R1 (DeepSeek-AI et al., 2025), QwQ, Qwen2.5 (Yang et al., 2024a), and Skywork-OR1 (He et al., 2025). See Supplementary Materials for details of evaluated LLMs. To reduce randomness, all presented results of models are average results over three trials.
Evaluation Methods. In the DRE-Bench benchmark, accuracy serves as the primary evaluation metric, defined as the proportion of samples for which the model’s output grid exactly matches the ground-truth output grid. To avoid contingency, each variable contains 12 samples for each value on average. All inferences are performed using the vLLM backend (Kwon et al., 2023). To ensure fairness and consistency, we adopt the official standardized prompting template released by ARCPrize (Prize, 2024).
@@
Average vs Human
Model-avg 57.01 59.21 23.50 46.57 57.15 33.04 8.53 32.91 23.72 3.63 31.78 19.71 0.80 5.72 0.00 2.17
Human-avg 75.56 82.22 68.89 75.56 91.11 75.56 46.67 71.11 73.33 24.44 88.89 62.22 46.67 77.78 17.78 47.41
4.2 Main Results in Four Levels
Furthermore, as task difficulty increases, performance disparities among models become more pronounced, highlighting the potential of incorporating dedicated reasoning paradigms for addressing fluid intelligence problems. For Level 4 tasks, which require conceptual knowledge, all existing models fail, underscoring the current limitations of even advanced reasoning models. These findings emphasize both the inherent challenges posed by our benchmark and its flexibility in revealing model capabilities across a wide spectrum of cognitive demands.
What’s more, we conduct a human study to validate our cognitive-aligned data framework. We extract 10% samples(about 400) from DRE-Bench based on its data distribution, and release a questionnaire to 20 annotators covering 10-40 age ranges. They are requested to fill out the test output as LLMs evaluated. We can observe in Table 1 that human accuracy also generally decreases as the level increases, which validates the justification of our 4-level framework. Compared with LLMs, human accuracy is slightly higher on average, indicating that existing LLMs have not yet reached human-level abstract reasoning, which is consistent with studies (Chollet, 2019; Chollet et al., 2025).
@@ 4.3 Dynamic Trends across Different Cognitive Levels
Since our generator is capable of producing data with varying levels of complexity, we conduct a fine-grained evaluation to assess model performance across data with different complexity. Figure 4 illustrates representative performance curves of nine LLMs for each cognitive level, with cases under the same rule gradually increasing in difficulty. More task curves are provided in Appendix E.3.
As Figure 4, since tasks on the Level-1 Attribute involve basic enumeration without substantial cognitive demands, most models consistently achieved high average accuracy, and increases in complexity had minimal impact. As for Level-2 Spatial, performance differences among models became increasingly pronounced, lower-performing models continued to struggle with even simple cases. Impressively, models with high accuracy remained robust, relatively unaffected by the increase in case complexity. This suggests that these models have, to some extent, acquired the capability to resolve spatial reasoning problems. Regarding tasks in Level-3 Sequential, we observe a substantial performance drop as the number of required planning steps increases. Most models can only manage the simplest scenarios, with a consistent failure point emerging when the planning depth reaches two steps. This highlights that current LLMs remain limited in intelligence and have yet to truly master such sequential rules. Finally, at Level-4 Conceptual, almost all models fail to provide correct solutions, even in the simplest cases under the gravity rule, indicating that current models have only a rudimentary grasp of physical concepts and have yet to internalize even the most fundamental principles of intuitive physics. In general, as task complexity increases across each cognitive level, the accuracy of models tends to decrease or fluctuate accordingly.
@@
To further illustrate the performance and stability of each model on dynamic task variants, Figure 5 presents the mean accuracy and corresponding variance across different cognitive levels. As shown in the figure, for the majority of Level-1 attribute tasks, OpenAI-o1, DeepSeek-R1, and Claude-3.7 demonstrate strong performance and high stability. However, when the task level increases to Level-2 spatial, Claude-3.7 exhibits substantial fluctuations in performance, indicating limited generalization capabilities at this level. In contrast, OpenAI-o1 and DeepSeek-R1 maintain comparable performance and stability to those observed at Level-1, highlighting the advantage of reasoning models in solving more cognitively demanding tasks. Moreover, in Level-3 sequential, most of the scatter points are concentrated in the lower-left region, suggesting that current models struggle to generalize effectively across the more complex and varied tasks at higher levels.
@@ 4.4 Ablation Study
Impact of the Number of In-context Learning Samples. Previous work (Brown et al., 2020; OpenAI, 2023) has demonstrated the effectiveness of in-context learning in enhancing model performance across LLM tasks. In this section, we investigate how the quantity of in-context samples affects performance in the abstract reasoning scenario. The results are shown in Figure 7. Overall, increasing the number of in-context samples helps models better capture underlying rules and improve performance. In higher levels like Level-2 Spatial, Level-3 Sequential and Level-4 Conceptual, increasing the number of in-context training samples leads to noticeable performance improvements. However, for Level-1 tasks, increasing the number of samples yields limited improvement. This suggests that adding more in-context examples has a limited impact when the model has already mastered the task or lacks the inherent capability to solve it.
Impact of the Auxiliary Visual Information. Previous studies (LeGris et al., 2024; Patterson et al., 2014) have shown that humans tend to perform better on abstract reasoning tasks when the grids are visualized, as visualization can aid in recognizing patterns and rules. Motivated by these findings, we investigate whether adding auxiliary visual information can enhance model performance. Specifically, we visualize each case by two formats: single-image, which presents all three training input-output pairs along with the test input in a single image; and multi-image, which provides them as seven separate images. Table 7 presents the experimental results of GPT-4o and Claude 3.7 across all four cognitive abstract reasoning levels. Overall, neither adding single-image nor multi-image format inputs can consistently outperform the text-only baseline, and in some instances, accuracy even declines. These results suggest that current models struggle to derive meaningful improvements in abstract reasoning from auxiliary visualized image inputs.
Figure 6: Accuracy of DeepSeek-R1 on different in-context training numbers.
Figure 7: Comparison of accuracy across text-only(-), single image(S-Img) and multi-image(M-Img) settings at four levels of cognitive reasoning.
@@
S-Img 96.84 17.14 31.25 15.87
M-Img 97.89 17.14 35.00 12.70
Impact of the Inference Time. It is demonstrated in  (DeepSeek-AI et al., 2025; OpenAI, 2024b; Qin et al., 2024; Huang et al., 2024) that inference-time scaling plays a crucial role in enhancing model performance on reasoning tasks. Building upon these, we take a step to examine how the model’s inference time varies as the complexity of reasoning tasks increases. According to related methods, we use the response latency to measure the inference time. The results are presented in Figure 9. We observe that at the low-level count task, as task complexity increases, the model tends to engage in deeper reasoning and can effectively maintain relatively stable and high accuracy. However, in high-level tasks (i.e., planning), even though the model’s inference time increases, it still fails to solve the more complex cases. This indicates that simply increasing inference time is insufficient to compensate for the model’s inherent limitations in high-level reasoning.
@@
Figure 9: Results of direction and symmetry.
@@ 4.5 Case Study
Analysis of Spatial Orientations. Upon closer examination of the results, we find that current models may demonstrate a distinct understanding of spatial orientation compared to humans. As shown in Table 9, the models achieve higher and more consistent accuracy in vertical (up/down) directions than in horizontal (left/right) ones in Move. Similarly, in symmetry tasks, performance is better for horizontal symmetry than for vertical symmetry. However, from the perspective of human cognition, directional distinctions are typically perceived as equivalent (Aflalo & Graziano, 2008; Ambinder et al., 2009). These findings suggest that current LLMs may exhibit systematic divergences from human cognitive patterns in processing spatial orientation.
@@
Analysis of Error Cases. As shown in Figure 10, we randomly select error cases from four cognitive levels and visualize the model output alongside the corresponding ground-truth for analysis. In Level-1 and Level-2, the differences between the model’s error predictions and the correct answers are relatively subtle, indicating that the model roughly understands the required operation. However, in Levels-3 and Level-4, the incorrect outputs become significantly more disorganized and divergent from the ground truth, suggesting a complete failure to grasp the underlying rule. This is especially evident in Level-4, where physical concepts pose substantial challenges to the models. These observations highlight that as the cognitive level increases, the nature of model errors becomes increasingly complex and unreasonable. The results of two auxiliary evaluation metrics: grid size precision and grid matching percentage in Appendix E.2 also confirm this circumstance.
@@
Figure 10: Error cases on o1: input, ground truth, and model output grids are visualized for each case.
5 Conclusion
@@
In this work, we present DRE-Bench, a benchmark designed to evaluate the fluid intelligence of large language models (LLMs) through abstract reasoning tasks. By combining a hierarchical task design, a scalable generator–solver pipeline, and dynamic task instantiation, DRE-Bench provides interpretability, scalability and robustness beyond prior benchmarks. Our experiments show that while reasoning-oriented models outperform general LLMs, their accuracy declines as cognitive level increases and case complexity rises. The results indicates that true fluid intelligence remains out of reach for current LLMs. DRE-Bench offers a principled framework for tracking reasoning progress and guide the development of future models with stronger generalizable intelligence.
Ethics Statement
This work complies fully with the ICLR Code of Ethics. No private, sensitive, or personally identifiable information was collected or used. The study involves no human subjects, no experiments on vulnerable populations, and no interventions requiring IRB approval. We confirm that our methodology and results do not raise foreseeable risks of harm, misuse, or ethical concerns beyond standard scientific research practices.
Reproducibility Statement
We present DRE-Bench, a benchmark for evaluating the fluid intelligence of large language models via abstract reasoning tasks structured in a four-level cognitive hierarchy. Compared with previous benchmarks, DRE-Bench probes latent rules across tasks and variants to provide interpretability, dynamic robustness, and scalability for tracking reasoning capabilities. We affirm the value of reproducibility in scientific research and therefore summarize the details of dataset, method, and experiments as follows:
  • Dataset. The detailed document and distribution of DRE-Bench are in Appendix C. And our dataset and all pairs of generator and solver have been available at the anonymous github link https://anonymous.4open.science/status/DRE-Bench-8098;
  • Method. The prompt templates to instruct code agent are detailed in Appendix D;
  • Experiment. Details about evaluated LLMs, results of two auxiliary evaluation metrics, more dynamic evaluation curves, example of two visual formats, and detailed table of variance are in Appendix E;
Appendix A Appendix
Appendix B LLM Usage Statement
We used LLMs(Gpt-5) to refine the writing, including checking grammar, polishing, and correcting typos. To ensure the writing quality, we further check and refine all the LLMs generated text. We assure that ideas, methods, code implementations, experiments, analyses, and conclusions aredone by human researchers ourselves.
@@ C.1 Detailed Dataset Content and Distribution
Symmetry Perform {Vertical, Horizontal, Center} symmetry of the object. number {1, 9} 75
Sequential Categorization Classify objects based on examples, and apply the corresponding rule to each category. category {1, 6} 65
@@
Conceptual Gravity Objects in mid-air should fall downward according to gravity. number {1, 9} 63
Reflection The light reflects upon hitting walls. number {1, 9} 100
Expansion Objects expand when heated until obstructed. number {1–9} 50
@@ C.2 Dataset Document
We provide comprehensive documentation of our dataset along with its intended use cases. The dataset and accompanying resources are available at the following link: https://anonymous.4open.science/status/DRE-Bench-8098, which includes metadata, format details, and so on.
Appendix D Details of Method
Our method employs two sequential system prompts to instruct code agent to implement generator and solver functions for each rule task. Based on the designed rule and corresponding constraints, the first system prompt guides the LLM to generate a structured code-like rule description, And the second system prompt translates this description into a complete Pygame program. We tested with different LLM-based code agents, including Gemini 2.5-Pro, Claude Opus 4-thinking, GPT-o3, and GPT-4o, and ultimately selected Gemini-2.5-Pro as the code agent in our experiments due to its higher success rate of generation. This two-prompt design ensures a clear division between rule modeling and executable code generation.
System Prompt 1: Rewrite the given rule and constraints into a structured rule description
You are an imaginative world architect and a technical artist. Your mission is to fuse a series of fundamental latent rules provided by the user (e.g., physics, math, artistic concepts) to create a concrete, detailed, and dynamic virtual scene.
Your output must adhere to the following guidelines:
1. **Structured Output**: Use a clear key-value format to describe the scene, making it easy to parse later.
2. **Code-like Description**: Use precise, quantifiable language, as if writing pseudocode or a configuration file. Avoid vague, literary descriptions.
3. **Dynamics and Interaction**: Focus on describing the behavior of elements, their interaction rules, and how they embody the users core rules.
Example Output Format:
Scene Name: [A creative name for the scene]
Core Rules: [Summarize the users concepts and how they are manifested in the scene]
Element List:
- Element A:
- Type: [e.g., Static Body, Dynamic Particle, Interactive Character]
- Visual Description: [A concise description of its appearance, material, color]
- Initial State: [Position coordinates, rotation angle, initial velocity, etc.]
- Behavioral Rules: [Describe how it moves, changes, and embodies the core concepts]
- Element B:
Physics & Interaction Rules:
- Rule 1: [e.g., Global gravity is set to a vector of (0, 0.1)]
- Rule 2: [e.g., When Element A and B collide, trigger a symmetrical bounce effect]
- Rule 3: [e.g., An element must find a path from a start to an end point, demonstrating pathfinding’]
System Prompt 2: Instruct the code agent to produce generator and solver fuctions based on the detailed rule description.
You are a senior Python game developer and an expert in using the Pygame library. Your task is to write a single, complete, and executable Pygame program that simulates the scene, strictly following the structured scene description provided by the user.
Your code must adhere to the following guidelines:
1. **Code Completeness**: Generate a single, complete Python script that includes all necessary Pygame initialization, the main loop, event handling, and rendering code.
2. **Precise Implementation**: The codes logic must accurately implement every element, behavior, and physical rule from the scene description.
3. **Readability**: The code must be clean and well-commented. Especially in the parts implementing core concepts (like gravity, pathfinding, rotation), explain how the code corresponds to the design document.
4. **No External Assets**: Use Pygames drawing functions (e.g., pygame.draw‘) to create geometric shapes. Do not rely on any external image or audio files.
Appendix E Experimental Details
E.1 Details of Evaluated LLMs